OpenStudy (anonymous):

What is the rate of change from x = 0 to x = ?

OpenStudy (anonymous):

@helper99 @amistre64

OpenStudy (freckles):

Not enough information given.

OpenStudy (anonymous):

x=0 to x=pie/2 sorry

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

So you aren't given a function?

OpenStudy (freckles):

You just want to know the change in x's?

OpenStudy (anonymous):

I'm given a graph but I don't know how to attach it.

OpenStudy (freckles):

Well yeah that is important information because with out a change of something over another change in something we do not have a rate of change

OpenStudy (anonymous):

How do I attach the graph?

OpenStudy (freckles):

anyways find the y value that corresponds to x=0 then also find the y value that corresponds to x=pi/2

OpenStudy (anonymous):

when x=0 y=0

OpenStudy (freckles):

ok and what y value corresponds to x=pi/2?

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

I know there's a formula for slope but Im not sure for this one

OpenStudy (freckles):

\[\frac{y_2-y_1}{x_2-x_1}\] will give you average rate of change

OpenStudy (anonymous):

ok let me calculate that

OpenStudy (anonymous):

sorry I can't figure this one out

OpenStudy (freckles):

so you are having trouble taking 0 from pi/2? And taking 0 from 3?

OpenStudy (anonymous):

it has to be -pie/6 pie/6 6/pie

OpenStudy (freckles):

You do know that a-0 is a right? or that 0-a=-a ?

OpenStudy (freckles):

Or are you having trouble with the division?

OpenStudy (freckles):

\[\frac{a}{\frac{b}{c}}=a \div \frac{b}{c}=a \cdot \frac{c}{b}\]

OpenStudy (anonymous):

how can i attach a graph?

OpenStudy (freckles):

if you simplify the fraction we have you will get one of the answers you mentioned above you can attach a file using the button below called "Attach File"

OpenStudy (freckles):

you should end up with a non fraction number on top and a fraction number on bottom but remember if you have a number over a fraction that is the same as saying that number times the reciprocal of that bottom fraction

OpenStudy (anonymous):

i got -pie/6 then

OpenStudy (freckles):

can you show me how you get that?

OpenStudy (freckles):

you are doing the change of y over the change of x right?

OpenStudy (freckles):

\[\frac{3-0}{\frac{\pi}{2}-0}\] since you said x=pi/2 corresponds to y=3 and x=0 corresponds to y=0

OpenStudy (anonymous):

I found the x and y coordinates that were found on the coresponding graph, and then I put those values into the slope formula

OpenStudy (freckles):

can you do 3-0 and pi/2-0?

OpenStudy (anonymous):

that would be 3/pie/2

OpenStudy (anonymous):

yikes i got a huge decimal number

OpenStudy (freckles):

the reason I ask is because somehow you got a negative but there is also another thing wrong with your answer you are having trouble dividing \[\frac{a}{\frac{b}{c}}=a \cdot \frac{c}{b}\]

OpenStudy (freckles):

and also pie isn't a number it is something you eat :p

OpenStudy (anonymous):

i got 3/2(pi)

OpenStudy (anonymous):

and yeah lol

OpenStudy (freckles):

you aren't going to try to do what I asked ?

OpenStudy (freckles):

\[\frac{a}{\frac{b}{c}}=a \cdot \frac{c}{b} \\ \\ \text{ you have } \\ \frac{3}{\frac{\pi}{2}} =3 \cdot \frac{2}{\pi}\]

OpenStudy (anonymous):

ok i did what you asked and i got 3.6366...

OpenStudy (freckles):

well if you leave it in exact form you can easily tell what your answer since your choices are in exact form that is none of your choices approximate the pi

OpenStudy (freckles):

also you shouldn't be getting 3.6366...

OpenStudy (freckles):

do you know how to multiply 3 and 2?

OpenStudy (anonymous):

6/pi?

OpenStudy (freckles):

finally I knew you knew how to multiply 3 and 2

OpenStudy (anonymous):

Thank you that was tough for me

OpenStudy (freckles):

i think you need to work on your division because it seems like that was most troubling for you

OpenStudy (anonymous):

I know how to use division, it's just that entire process was confusing me

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (freckles):

in general this is how you divide a integer by a rational: ( that an integer by a fraction) \[\frac{a}{\frac{b}{c}}=\] \[\frac{a}{\frac{b}{c}} \cdot \frac{\frac{c}{b}}{\frac{c}{b}}=\frac{a \cdot \frac{c}{b}}{\frac{b}{c} \cdot \frac{c}{b} } =\frac{ a \cdot \frac{c}{b}}{(1)}=a \cdot \frac{c}{b}=\frac{a c}{b}\]

OpenStudy (freckles):

like to do a/(b/c) you just do a *c/b

OpenStudy (anonymous):

yeah that seems to be the best way

OpenStudy (freckles):

that is the only way I know how to divide integers by fractions :p

OpenStudy (freckles):

do you know another way?

OpenStudy (freckles):

anyways I don't there is another way

OpenStudy (freckles):

lol but what is your other question

OpenStudy (anonymous):

What sine function represents an amplitude of 4, a period of , no horizontal shift, and a vertical shift of −3?

OpenStudy (anonymous):

this would be best if there was a way that i could attach the answers

OpenStudy (freckles):

well there is

OpenStudy (freckles):

you can attach a file using the button below

OpenStudy (anonymous):

i know but how do i transfer the info to a file?

OpenStudy (freckles):

by anyways \[y=A \sin(B(x-C))+D\] The amp is |A| horizontal shift is C Vertical shift is D The period is 2pi/B

OpenStudy (freckles):

Anyways let's pretend the period is k since we aren't given period what would that make B in terms of k? well \[\frac{2 \pi}{B}=k \] we would just need to solve this for B \[2 \pi=k B \text{ After multiplying } B \text{ on both sides } \\ \frac{ 2\pi}{k}=B \text{ after dividing both sides by } k\] so putting this in so far we have: \[y=A \sin(B(x-C))+D \\ y=A \sin(\frac{2 \pi}{k}(x-C))+D\]

OpenStudy (anonymous):

I think it would be f(x)equals 4sin(pi/2-3)

OpenStudy (freckles):

Now where they say there is no horizontal shift that means C value is?

OpenStudy (anonymous):

0?

OpenStudy (freckles):

yes so replace C with 0

OpenStudy (freckles):

so you have so far \[y=A \sin(B(x-C))+D \\ y=A \sin(\frac{2 \pi}{k}(x-0))+D \text{ or } y=A \sin(\frac{2 \pi}{k}x)+D\]

OpenStudy (anonymous):

I think it would be f(x)equals 4sin(pi/2-3)

OpenStudy (freckles):

now D is the vertical shift number and you are given the vertical shift is -3 so D is ?

OpenStudy (freckles):

\[y=A \sin(B(x-C))+D \\ y=A \sin(\frac{2 \pi}{k}(x-0))+D \text{ or } y=A \sin(\frac{2 \pi}{k}x)+D \\ y=4 \sin(\frac{2 \pi}{k} x)-3 \text{ or it could be } y=-4 \sin(\frac{2 \pi}{k} x)-3\] since you know it didn't say if it has parent function y=sin(x) or if it had y=-sin(x) like the amp=4 could be mean that result came from |4| or |-4|

OpenStudy (freckles):

anyways either one of those should right given the description above we still don't know what k was because we were never given the period